Marquis:

Evaluate the expression 2003-02-05-00-00_files/i0290000.jpg for x = 11 and y = –7.

Eiwoh2:

We cannot see those files

Eiwoh2:

DO SOMETHING ABOUT THIS, ULTRI!!!

Ultrilliam:

if its formatted like that

Ultrilliam:

\[\large\left(\begin{matrix} \frac{x^{-3}y^7}{x^3y^6} \end{matrix}\right)^0\]